OpenStudy (anonymous):

What is 15 degrees Celsius (°C) in degrees Fahrenheit (°F)? Use the formula F = 1.8C + 32. A. 20.5 °F B. 37 °F C. 59 °F D. 84.6 °F

OpenStudy (charlotte123):

the answer would be c :D

OpenStudy (anonymous):

\[F = 1.8C + 32.\]and you are told \(C=15\)\[F = 1.8\times 15+ 32.\]

OpenStudy (charlotte123):

1.8(15) + 32 = 59 :D